OpenStudy (anonymous):

Simplify the expression 5(x + 3) - 6(5x + 9).

mathslover (mathslover):

\[\huge{5x+15-30x-54}\] \[\huge{-25x-39}\] remember : \[\huge{a(b+c)=ab+ac}\]

mathslover (mathslover):

u must know how to combine like terms like here : \[\large{\color{blue}{5x+15-30x-54=5x-30x+15-54}}\]

OpenStudy (anonymous):

Thanks @mathslover your ALOT of help :)